# SiT: Exploring Flow and Diffusion-based Generative Models with Scalable Interpolant Transformers

### v1

@@ Abstract
We present Scalable Interpolant Transformers (SiT), a family of generative models built on the backbone of Diffusion Transformers (DiT). The interpolant framework, which allows for connecting two distributions in a more flexible way than standard diffusion models, makes possible a modular study of various design choices impacting generative models built on dynamical transport: using discrete vs. continuous time learning, deciding the objective for the model to learn, choosing the interpolant connecting the distributions, and deploying a deterministic or stochastic sampler. By carefully introducing the above ingredients, SiT surpasses DiT uniformly across model sizes on the conditional ImageNet 256x256 benchmark using the exact same backbone, number of parameters, and GFLOPs. By exploring various diffusion coefficients, which can be tuned separately from learning, SiT achieves an FID-50K score of 2.06.
@@ 1 Introduction
Contemporary success in image generation has come from a combination of algorithmic advances and improvements in model architecture and progress in scaling neural network models and data. State-of-the-art diffusion models [25, 51] proceed by incrementally transforming data into Gaussian noise as prescribed by an iterative stochastic process, which can be specified either in discrete or continuous time. At an abstract level, this corruption process can be viewed as defining a time-dependent distribution that is iteratively smoothed from the original data distribution into a standard normal distribution. Diffusion models learn to reverse this corruption process and push Gaussian noise backwards along this connection to obtain data samples. The objects learned to perform this transformation are conventionally either predicting the noise in the corruption process [25] or predicting the score of the distribution connecting the data and the Gaussian [62], though alternatives of these choices exist [54, 27].
@@
Table 1: Scalable Interpolant Transformers. We systematically vary the following aspects of a generative model: discrete vs. continuous time, model prediction, interpolant, choice of sampler. The resulting Scalable Interpolant Transformer (SiT) model, under identical training compute, consistently outperforms the Diffusion Transformer (DiT) in generating 256×256 ImageNet images. All models employ a patch size of 2. In this work, we ask the question: What is the source of the performance gain?
@@
The neural network architectures used to represent these objects have been shown to perform well on a variety of tasks. While diffusion models were originally built upon a U-Net backbone [25, 52], recent work has highlighted that architectural advances in vision such as the Vision Transformer (ViT) [21] can be incorporated into the standard diffusion model pipeline to improve performance [48]. The aims of [48] were to push improvements on the model side of the duality of algorithm and model.
Orthogonally, significant research effort has gone into exploring the structure of the noising process, which has been shown to lead to performance benefits [36, 32, 35, 58]. Yet, many of these efforts do not move past the notion of passing data through a diffusion process with an equilibrium distribution, which is a restricted type of connection between the data and the Gaussian. The recently-introduced stochastic interpolants [3] lift such constraints and introduce more flexibility in the noise-data connection. In this paper, we further explore its performance in large scale image generation.
Intuitively, we expect that the difficulty of the learning problem can be related to both the specific connection chosen and the object that is learned. Our aim is to clarify these design choices, in order to simplify the learning problem and improve performance. To glean where potential benefits arise in the learning problem, we start with Denoising Diffusion Probabilistic Models (DDPMs) and sweep through adaptations of: (i) which object to learn, and (ii) which interpolant to choose to reveal best practices.
@@
Figure 2: SiT observes improvement in FID across all model sizes. We show FID-50K over training iterations for both DiT and SiT. All results are produced by a Euler-Maruyama sampler using 250 integration steps. Across all model sizes, SiT converges much faster.
  • By moving from discrete to continuous time, changing the model prediction, interpolant, and the choice of sampler, we observe a consistent performance improvement over the Diffusion Transformer (DiT).
  • We systematically study where these improvements come from by addressing these factors one by one: learning in continuous time; learning a velocity as compared to a score; changing the interpolant connecting the the two distributions; and using the velocity in an SDE sampler with particular choices of diffusion coefficients.
  • We show that the SDE for the interpolant can be instantiated using just a velocity model, which we use to push the performance of these methods beyond previous results.
@@ 2.1 Flows and diffusions
In recent years, a flexible class of generative models based on turning noise ${\boldsymbol{\varepsilon}}\sim\mathsf{N}(0,\mathbf{I})$ into data $\mathbf{x}_{*}\sim p(\mathbf{x})$ have been introduced. These models use the time-dependent process
@@
where $\alpha_{t}$ is a decreasing function of $t$ and $\sigma_{t}$ is an increasing function of $t$. Stochastic interpolants and other flow matching methods [1, 3, 39, 42] restrict the process (1) on $t\in[0,1]$, and set $\alpha_{0}=\sigma_{1}=1$, $\alpha_{1}=\sigma_{0}=0$, so that ${\mathbf{x}}_{t}$ interpolates exactly between ${\mathbf{x}}_{*}$ at time $t=0$ and ${\boldsymbol{\varepsilon}}$ and time $t=1$. By contrast, score-based diffusion models [62, 36, 32] set both $\alpha_{t}$ and $\sigma_{t}$ indirectly through different formulations of a stochastic differential equation (SDE) with $\mathsf{N}(0,\mathbf{I})$ as its equilibrium distribution. Moreover, they consider the process ${\mathbf{x}}_{t}$ on an interval $[0,T]$ with $T$ large enough that ${\mathbf{x}}_{T}$ approximates a Gaussian distribution.
Probability flow.
Common to both stochastic interpolants and score-based diffusion models is the observation that the process ${\mathbf{x}}_{t}$ can be sampled dynamically using either an SDE or a probability flow ordinary differential equation (ODE). More precisely, the marginal probability distribution $p_{t}({\mathbf{x}})$ of ${\mathbf{x}}_{t}$ in (1) coincides with the distribution of the probability flow ODE with a velocity field
@@
Equation (3) is derived in Sec. A.1. By solving the probability flow ODE (2) backwards in time from ${\mathbf{X}}_{T}={\boldsymbol{\varepsilon}}\sim\mathsf{N}(0,\mathbf{I})$, we can generate samples from $p_{0}({\mathbf{x}})$, which approximates the ground-truth data distribution $p({\mathbf{x}})$. We refer to (2) as a flow-based generative model.
@@
$d{\mathbf{X}}_{t}={\mathbf{v}}({\mathbf{X}}_{t},t)dt+\frac{1}{2}w_{t}{\mathbf{s}}({\mathbf{X}}_{t},t)dt+\sqrt{w_{t}}d\bar{\mathbf{W}}_{t},$ (4)
where $\bar{\mathbf{W}}_{t}$ is a reverse-time Wiener process, $w_{t}>0$ is an arbitrary time-dependent diffusion coefficient, ${\mathbf{v}}({\mathbf{x}},t)$ is the velocity defined in (3), and where ${\mathbf{s}}({\mathbf{x}},t)=\nabla\log p_{t}({\mathbf{x}})$ is the score. Similar to ${\mathbf{v}}$, this score is given by the conditional expectation
@@
This equation is derived in Sec. A.3. Similarly, solving the reverse SDE (4) backwards in time from ${\mathbf{X}}_{T}={\boldsymbol{\varepsilon}}\sim\mathsf{N}(0,\mathbf{I})$ enables generating samples from the approximated data distribution $p_{0}(\mathbf{x})\sim p({\mathbf{x}})$. We refer to (2) as a diffusion-based generative model.
@@ 2.2 Estimating the score and the velocity
We note that any time-dependent weight can be included under the integrals in both (6) and (7). These weight factors are key in the context of score-based models when $T$ becomes large [35]; in contrast, with stochastic interpolants where $T=1$ without any bias, these weights are less important and might impose numerical stability issue (see Appendix B).
@@
We observed that only one of the two quantities ${\mathbf{s}}_{\theta}({\mathbf{x}},t)$ and ${\mathbf{v}}_{\theta}({\mathbf{x}},t)$ needs to be estimated in practice. This follows directly from the constraint
@@
We will use this relation to specify our model prediction. Conversely, we can also express ${\mathbf{v}}({\mathbf{x}},t)$ in terms of ${\mathbf{s}}({\mathbf{x}},t)$. In our experiments, we typically learn the velocity field ${\mathbf{v}}({\mathbf{x}},t)$ and use it to express the score ${\mathbf{s}}({\mathbf{x}},t)$ when using an SDE for sampling. We include a detailed derivation in Sec. A.4.
Note that by our definitions $\dot{\alpha}_{t}<0$ and $\dot{\sigma}_{t}>0$, so that the denominator of (9) is never zero. Yet, $\sigma_{t}$ vanishes at $t=0$, making the $\sigma_{t}^{-1}$ in (9) appear to cause a singularity there [^1] We remark that ${\mathbf{s}}({\mathbf{x}},t)$ can be shown to be non-singular at $t=0$ analytically if the data distribution $p({\mathbf{x}})$ has a smooth density [3], though this singularity appears in numerical implementations and losses in general.. This suggests the choice $w_{t}=\sigma_{t}$ in (4) to cancel this singularity (see Sec. A.3), for which we will explore the performance in the numerical experiments.
@@ 2.3 Specifying the interpolating process
In Score-Based Diffusion Models (SBDM), the choice of $\alpha_{t}$ and $\sigma_{t}$ in (1) is typically determined by the choice of the forward SDE used to define this process, though recent work has tried to reconsider this [32, 35]. For example, if we use the standard variance-preserving (VP) SDE [62]
@@
for some $\beta_{t}>0$, it can be shown (see Appendix B) that the solution to (10) has the same probability distribution $p_{t}({\mathbf{x}})$ as the process ${\mathbf{x}}_{t}$ defined in (1) for the choice
@@
The only design flexibility in (11) comes from the choice of $\beta_{t}$, because it determines both $\alpha_{t}$ and $\sigma_{t}$222 VP is the only linear scalar SDE with an equilibrium distribution [58]; interpolants extend beyond $\alpha_{t}^{2}+\sigma_{t}^{2}=1$ by foregoing the requirement of an equilibrium distribution. . For example, setting $\beta_{t}=1$ leads to $\alpha_{t}=e^{-t}$ and $\sigma_{t}=\sqrt{1-e^{-2t}}$. This choice necessitates taking $T$ sufficiently large [25] or searching for more appropriate choices of $\beta_{t}$ [62, 16, 58] to reduce the bias induced by the fact that the solution to the SDE (10) only converges to ${\boldsymbol{\varepsilon}}\sim\mathsf{N}(0,\mathbf{I})$ as $t\to\infty$.
@@
  1. (i)
    $\alpha_{t}^{2}+\sigma_{t}^{2}>0$ for all $t\in[0,1]$;
  2. (ii)
    $\alpha_{t}$ and $\sigma_{t}$ are differentiable for all $t\in[0,1]$;
  3. (iii)
    $\alpha_{1}=\sigma_{0}=0$, $\alpha_{0}=\sigma_{1}=1$;
@@ 2.4 Specifying the diffusion coefficient
As stated earlier, the SBDM diffusion coefficient used in the reverse SDE (4) is usually taken to match that of the forward SDE (10). That is, one sets $w_{t}=\beta_{t}$. In the stochastic interpolant framework, this choice is again subject to greater flexibility: any $w_{t}>0$ can be used. Interestingly, this choice can be made after learning, as it does not affect the velocity ${\mathbf{v}}({\mathbf{x}},t)$ or the score ${\mathbf{s}}({\mathbf{x}},t)$. In our experiments, we exploit this flexibility by considering the choices listed in Table 2.
Expression for $w_{t}$
$\beta_{t}=-2\sigma_{t}(\dot{\sigma}_{t}-\frac{\sigma_{t}\dot{\alpha}_{t}}{\alpha_{t}})$
$\sigma_{t}$
$1-t$
$\sin^{2}(\pi t)$
$(\cos(\pi t)+1)^{2}$
$(\cos(\pi t)-1)^{2}$
Table 2: Diffusion coefficients. They can be specified after learning to maximize performance. $w_{t}=\beta_{t}$ in the first row corresponds to SBDM (see Eq. 11), which is coupled to the forward process. A detailed derivation is provided in Appendix B. $w_{t}=\sigma_{t}$ in the second row is used to eliminate the singularity at $t=0$ following the explanation at the end of Sec. 2.2. In other choices we decouple $w_{t}$ from the interpolant and experiment with removing diffusivity at different times in sampling.
2.5 Time-discretization and link with DDPM
During inference, continuous time models must be discretized when solving the probability flow ODE (2) and the reverse-time SDE (4). This allows us to make a link with DDPMs [25].
Assuming that we discretize time using a grid $0=t_{0}<t_{1}<t_{2}<\ldots<t_{N}=T$, the process (1) can be evaluated at each grid point, ${\mathbf{x}}_{t_{i}}=\alpha_{t_{i}}{\mathbf{x}}_{*}+\sigma_{t_{i}}{\boldsymbol{\varepsilon}}$, and both the velocity and the score can be estimated on these points via the losses
$\mathcal{L}^{N}_{\mathrm{s}}(\theta)$ $=\sum_{i=0}^{N}\mathbb{E}[\|\sigma_{t_{i}}{\mathbf{s}}_{\theta}({\mathbf{x}}_{t_{i}},{t_{i}})+{\boldsymbol{\varepsilon}}\|^{2}],$ (13)
$\mathcal{L}^{N}_{\mathrm{v}}(\theta)$ $=\sum_{i=0}^{N}\mathbb{E}[\|{\mathbf{v}}_{\theta}({\mathbf{x}}_{t_{i}},{t_{i}})-\dot{\alpha}_{t_{i}}{\mathbf{x}}_{*}-\dot{\sigma}_{t_{i}}{\boldsymbol{\varepsilon}}\|^{2}].$ (14)
Moreover, only the learned ${\mathbf{s}}_{\theta}({\mathbf{x}},{t_{i}})$ or ${\mathbf{v}}_{\theta}({\mathbf{x}},{t_{i}})$ is needed to integrate the probability flow ODE (2) and the reverse-time SDE (4) on the same grid. The resulting procedure, in which we define ${\mathbf{x}}_{t}$ iteratively on the grid, is a generalization of DDPM. Starting from ${\mathbf{x}}_{t_{0}}={\mathbf{x}}_{*}$, we set for $i\geq 0$,
${\mathbf{x}}_{t_{i+1}}$ $=\sqrt{1-h\beta_{t_{i}}}{\mathbf{x}}_{t_{i}}+\sqrt{h\beta_{t_{i}}}{\boldsymbol{\varepsilon}}_{t_{i}},$ (15)
where $h=t_{i+1}-t_{i}$ and where we assume that the grid is uniform. Because $\sqrt{1-h\beta_{t_{i}}}=1-\frac{1}{2}h\beta_{t_{i}}+o(h)$, it is easy to see that (15) is a consistent time-discretization of the forward SDE (10). Our results show that it is not necessary to specify the time discretized process ${\mathbf{x}}_{t_{i}}$ using (15), but instead we can directly use (1) on the time grid.

### v2

@@ Abstract
We present Scalable Interpolant Transformers (SiT), a family of generative models built on the backbone of Diffusion Transformers (DiT). The interpolant framework, which allows for connecting two distributions in a more flexible way than standard diffusion models, makes possible a modular study of various design choices impacting generative models built on dynamical transport: learning in discrete or continuous time, the objective function, the interpolant that connects the distributions, and deterministic or stochastic sampling. By carefully introducing the above ingredients, SiT surpasses DiT uniformly across model sizes on the conditional ImageNet $256\times 256$ and $512\times 512$ benchmark using the exact same model structure, number of parameters, and GFLOPs. By exploring various diffusion coefficients, which can be tuned separately from learning, SiT achieves an FID-50K score of 2.06 and 2.62, respectively. Code is available here: https://github.com/willisma/SiT
@@ 1 Introduction
Contemporary success in image generation has come from a combination of algorithmic advances, improvements in model architecture, and progress in scaling neural network models and data. State-of-the-art diffusion models [25, 53] proceed by incrementally transforming data into Gaussian noise as prescribed by an iterative stochastic process, which can be specified either in discrete or continuous time. At an abstract level, this corruption process can be viewed as defining a time-dependent distribution that is iteratively smoothed from the original data distribution into a standard normal distribution. Diffusion models learn to reverse this corruption process and push Gaussian noise backwards along this connection to obtain data samples. The objects learned to perform this transformation conventionally predict either the noise in the corruption process [25] or the score of the distribution that connects the data and the Gaussian [64], though alternatives of these choices exist [56, 28].
Table 1: Scalable Interpolant Transformers. We systematically vary the following aspects of a generative model: time discretization, model prediction, interpolant, and sampler. The resulting Scalable Interpolant Transformer (SiT) model, under identical training compute, consistently outperforms the Diffusion Transformer (DiT) in generating 256×256 ImageNet images. All models employ a patch size of 2. In this work, we ask the question: What is the source of the performance gain?
@@
Figure 2: SiT improves FID across all model sizes. FID-50K over training iterations for both DiT and SiT. All results are produced by a Euler-Maruyama sampler using 250 integration steps. Across all model sizes, SiT converges much faster.
While diffusion models originally represented these objects with a U-Net architecture [25, 54], recent work has highlighted that architectural advances in vision such as the Vision Transformer (ViT) [21] can be incorporated into the standard diffusion model pipeline to improve performance [50].
Orthogonally, significant research effort has gone into exploring the structure of the noising process, which has been shown to lead to performance benefits [37, 33, 36, 60]. Yet, many of these efforts do not move past the notion of passing data through a diffusion process with an equilibrium distribution, which is a restricted type of connection between the data and the Gaussian. Recently-introduced stochastic interpolants [2] lift this constraint and introduce more flexibility in the noise-data connection. In this paper, we systematically explore the effect of this flexibility on performance in large scale image generation.
Intuitively, we expect that the difficulty of the learning problem can be related to both the specific connection chosen and the object that is learned. Our aim is to clarify these design choices, so as to simplify the learning problem and thereby improve performance. To understand where potential benefits arise in the learning problem, we start with Denoising Diffusion Probabilistic Models (DDPMs) and sweep through adaptations of: (i) which object to learn, and (ii) which interpolant to choose to reveal best practices.
@@
  • We systematically study the SiT design space through the combinations of the four key components: time discretization, model prediction, interpolant, and sampler.
  • We provide theoretical motivation for the choice of each component and study how they lead to improved practical performance.
  • We exploit the tunability of the diffusion coefficient of the stochastic sampler, and show that its adaptation can tighten control of the KL-divergence between the model and the target. We show how this leads to empirical benefits without any additional re-training.
  • Combining the best design choices identified in each component, our SiT model surpasses Diffusion Transformer(DiT) on both $256\times 256$ and $512\times 512$ image resolution, achieving FID-50K scores of 2.06 and 2.62, respectively, without modifying any structure or hyperparameter of the model.
@@ 2.1 Flows and diffusions
Flow and diffusion models both utilize stochastic processes to gradually turn noise ${\boldsymbol{\varepsilon}}\sim\mathsf{N}(0,\mathbf{I})$ into data $\mathbf{x}_{*}\sim p(\mathbf{x})$ for the generating task. Such time-dependent processes can be summarized as follow
@@
where $\alpha_{t}$ is a decreasing function of $t$ and $\sigma_{t}$ is an increasing function of $t$. Stochastic interpolants and other flow matching methods [4, 2, 41, 43] restrict the process (1) on $t\in[0,1]$, and set $\alpha_{0}=\sigma_{1}=1$, $\alpha_{1}=\sigma_{0}=0$, so that ${\mathbf{x}}_{t}$ interpolates exactly between ${\mathbf{x}}_{*}$ at time $t=0$ and ${\boldsymbol{\varepsilon}}$ and time $t=1$. By contrast, score-based diffusion models [64, 37, 33] set both $\alpha_{t}$ and $\sigma_{t}$ indirectly through a forward-time stochastic differential equation (SDE) with $\mathsf{N}(0,\mathbf{I})$ as its equilibrium distribution, i.e. ${\mathbf{x}}_{t}$ converges to $\mathsf{N}(0,\mathbf{I})$ only if $t\to\infty$.
Despite the nuances in formulating the stochastic processes ${\mathbf{x}}_{t}$, common to both stochastic interpolants and score-based diffusion models is the observation that ${\mathbf{x}}_{t}$ can be sampled dynamically using either a reverse-time SDE or a probability flow ordinary differential equation (ODE).
Probability flow ODE.
The marginal probability distribution $p_{t}({\mathbf{x}})$ of ${\mathbf{x}}_{t}$ in (1) coincides with the distribution of the probability flow ODE with a velocity field
@@
The correspondence between $p_{t}({\mathbf{x}})$ and (2) and the formulation of (3) is derived in Appendix 0.A.1. By solving (2) backwards in time from ${\mathbf{X}}_{T}={\boldsymbol{\varepsilon}}\sim\mathsf{N}(0,\mathbf{I})$, we can generate samples from $p_{0}({\mathbf{x}})$, which approximates the ground-truth data distribution $p({\mathbf{x}})$. We refer to (2) as a flow-based generative model.
@@
${\mathrm{d}}{\mathbf{X}}_{t}={\mathbf{v}}({\mathbf{X}}_{t},t){\mathrm{d}}t-\frac{1}{2}w_{t}{\mathbf{s}}({\mathbf{X}}_{t},t){\mathrm{d}}t+\sqrt{w_{t}}{\mathrm{d}}\bar{\mathbf{W}}_{t},$ (4)
where $\bar{\mathbf{W}}_{t}$ is a reverse-time Wiener process, $w_{t}>0$ is an arbitrary time-dependent diffusion coefficient, ${\mathbf{v}}({\mathbf{x}},t)$ is the velocity defined in (3), and ${\mathbf{s}}({\mathbf{x}},t)=\nabla\log p_{t}({\mathbf{x}})$ is the score. Similar to ${\mathbf{v}}$, this score is given by the conditional expectation
@@
Again, the correspondence between $p_{t}({\mathbf{x}})$ and (4) and the formulation of (5) is derived in Appendix 0.A.3. Solving the reverse SDE (4) backwards in time from ${\mathbf{X}}_{T}={\boldsymbol{\varepsilon}}\sim\mathsf{N}(0,\mathbf{I})$ enables generating samples from the approximated data distribution $p_{0}(\mathbf{x})\sim p({\mathbf{x}})$. We refer to (4) as a stochastic generative model.
@@ 2.2 Estimating the score and the velocity
We note that any time-dependent weight can be included under the integrals in both (6) and (7). These weight factors are key in the context of score-based models when $T$ becomes large [36]; in contrast, with stochastic interpolants where $T=1$ without any bias, these weights are less important and might impose numerical stability issue (see Appendix 0.B).
@@
We observed that only one of ${\mathbf{s}}_{\theta}({\mathbf{x}},t)$ and ${\mathbf{v}}_{\theta}({\mathbf{x}},t)$ is needed to be estimated in practice. This follows directly from the constraint
@@
We include a detailed derivation in Appendix 0.A.4. Notably, given the simply linear relationship posed by (9), we can also express ${\mathbf{v}}({\mathbf{x}},t)$ in terms of ${\mathbf{s}}({\mathbf{x}},t)$. We will use this relation to specify our model prediction. In our experiments, we typically learn the velocity field ${\mathbf{v}}({\mathbf{x}},t)$ and use it to express the score ${\mathbf{s}}({\mathbf{x}},t)$ when using an SDE for sampling.
Note that by our definitions $\dot{\alpha}_{t}<0$ and $\dot{\sigma}_{t}>0$, so that the denominator of (9) is never zero. Yet, $\sigma_{t}$ vanishes at $t=0$, making the $\sigma_{t}^{-1}$ in (9) cause a singularity [^1] We remark that ${\mathbf{s}}({\mathbf{x}},t)$ can be shown to be non-singular at $t=0$ analytically if the data distribution $p({\mathbf{x}})$ has a smooth density [2], though this singularity appears in numerical implementations and losses in general.. This suggests the choice $w_{t}=\sigma_{t}$ in (4) to cancel this singularity, for which we will explore the performance in the numerical experiments.
Time discretization.
The objective functions specified above are defined over a continuous time domain, as opposed to DDPM which couples the time grid used in learning to that used in sampling. Learning in continuous time allows us to specify a discretization used in sampling a posteriori, which allows for flexibility in both sampling efficiency and performance.
@@ 2.3 Specifying the interpolating process
In Sec. 2.1 we present the general definition of interpolants ($\alpha_{t}$ and $\sigma_{t}$) for both stochastic interpolant and score-based diffusion. In this section we dive into more details and specify the three choices of interpolants to explore in the experiments.
@@
We follow [64] and use the standard variance-preserving (VP) SDE in forward-time
@@
for some $\beta_{t}>0$, ${\mathbf{x}}_{t}$’s perturbation kernel $p_{t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})=\mathsf{N}(\alpha_{t}{\mathbf{x}}_{t},\sigma_{t}^{2}\mathbf{I})$ is defined by
@@
The only design flexibility in (11) comes from the choice of $\beta_{t}$, as it determines both $\alpha_{t}$ and $\sigma_{t}$222 VP is the only linear scalar SDE with an equilibrium distribution [60]; interpolants extend beyond $\alpha_{t}^{2}+\sigma_{t}^{2}=1$ by foregoing the requirement of an equilibrium distribution. . For example, setting $\beta_{t}=1$ leads to $\alpha_{t}=e^{-t}$ and $\sigma_{t}=\sqrt{1-e^{-2t}}$. This choice necessitates taking $T$ sufficiently large [25] or searching for more appropriate choices of $\beta_{t}$ [64, 16, 60] to reduce the bias. To be specific, such bias comes from the mismatch between the condition $\varepsilon\sim\mathsf{N}(0,\mathbf{I})$ used in practice for sampling and the density of the process ${\mathbf{x}}_{1}\not\sim\mathsf{N}(0,\mathbf{I})$, as stated in Sec. 2.1.
@@
(i) $\alpha_{t}^{2}+\sigma_{t}^{2}>0$ for all $t\in[0,1]$;
(ii) $\alpha_{t}$ and $\sigma_{t}$ are differentiable for all $t\in[0,1]$;
(iii) $\alpha_{1}=\sigma_{0}=0$, $\alpha_{0}=\sigma_{1}=1$;
@@ 2.4 Specifying the diffusion coefficient
As stated earlier, the SBDM diffusion coefficient used in (4) is usually taken to match that of (10). That is, one sets $w_{t}=\beta_{t}$. In the stochastic interpolant framework, this choice is again subject to greater flexibility: any $w_{t}\geq 0$ can be used. Interestingly, this choice can be made after learning, as it does not affect the velocity ${\mathbf{v}}({\mathbf{x}},t)$ or the score ${\mathbf{s}}({\mathbf{x}},t)$. In our experiments, we exploit this flexibility by considering the following choices:
  1. (i)
    $w_{t}=\sigma_{t}$; this is used to eliminate the singularity at $t=0$ following the explanation at the end of Sec. 2.2;
  2. (ii)
    $w_{t}=\sin^{2}(\pi t)$; this also eliminates the singularity at $t=0$, and allows us to explore the effect of removing diffusivity at times close to $t=1$ in sampling.
  3. (iii)
    $w_{t}$ can be chosen to minimize an upper bound on the KL divergence $D_{\mathrm{KL}}(p({\mathbf{x}})\|p_{0}({\mathbf{x}}))$, where $p({\mathbf{x}})$ denotes the true data distribution and $p_{0}({\mathbf{x}})$ refers to the density of ${\mathbf{x}}_{t}$ at $t=0$. Disregarding the simulation cost of integrating the SDE (4), it can be shown (see Appendix 0.A.5) that the following choice of $w_{t}$ minimizes the KL upper bound:
    $w_{t}=w_{t}^{\mathrm{KL}}\equiv 2\left(\dot{\sigma}_{t}\sigma_{t}-\frac{\dot{\alpha}_{t}\sigma_{t}^{2}}{\alpha_{t}}\right).$ (13)
    Under the SBDM-VP interpolant, $w^{\mathrm{KL}}_{t}$ coincides with $\beta_{t}$; this aligns with the claim made in [63].
  4. (iv)
    If the SDE in (iii) becomes hard to integrate because of the magnitude of $w_{t}^{\mathrm{KL}}$ near $t=1$, one may wish to regularize the diffusion coefficient to reduce the integration cost. For example, difficulties may arise for the Linear and GVP interpolants, because $w_{t}^{\mathrm{KL}}\to\infty$ as $t\to 1$ given the presence of $\alpha_{t}$ in the denominator of (13). Including the integration cost of (4), it can also be shown (see Appendix 0.A.5) that an optimal regularized $w_{t}$ is given by
    $w_{t}^{\mathrm{KL},\eta}\equiv w_{t}^{\mathrm{KL}}\sqrt{\frac{\mathcal{L}_{t}}{\mathcal{L}_{t}+2\eta(w_{t}^{\mathrm{KL}})^{2}}},$ (14)
    where $\mathcal{L}_{t}$ is the value of $\mathcal{L}_{\mathrm{v}}$ in Sec. 2.2 at time $t$, and $\eta$ is any non-negative constant. With $\eta=0$, we recover $w_{t}^{\mathrm{KL}}$. For score models, we first convert to a velocity model following (9), then calculate the corresponding $\mathcal{L}_{\mathrm{v}}$. As $t\to 1$, $w_{t}^{\mathrm{KL},\eta}$ approaches a limit at $\sqrt{\frac{\mathcal{L}_{t\to 1}}{2\eta}}$. If $\mathcal{L}_{t}$ is defined everywhere on $[0,1]$, then $w_{t}^{\mathrm{KL},\eta}$ will be well-behaved on $[0,1]$.